Recent Maintenance – ESK

ESK operating status: up

2024.03.04

Calibration results show STS-1N long-period high pass filter failed. Trip to replace STS-1 sensors planned for Spring 2024.

2015.01.12

An IDA engineer is on-site to relocate the station computers and ancillary equipment to the vault, and upgrade the power system.

2014.11.19

The station was down on day 2014-321-09:58 due to a power problem.  IDA engineers instructed the station operator to bypass the UPS and data acquisition was restored.

2012.03.20

IDA station ESK is back on-line as of day 2012,080-05:26 after the station operator restarted the fibre-optic transciever and cycled power to the digitizers.

2012.03.19

Further communication with BGS (British Geological Survey) indicates there was a power outage around 2012,066-17:11 which caused loss of communication between the station computer and Q330 digitizers. Due to the digitizers’ limited data buffer the station is down and IDA engieers continue to work with the station operator to resolve this problem.

2012.03.12

Telemetry is currently down for IDA station ESK. However, the station is still acquiring data. The station operator has been requested to investigate.

2010.02.25

An IDA engineer visited the station and upgraded it to a Quanterra Q330HR system. Seismometers include an STS-2 for broadband and short period channels, and an episensor strong motion detector.

2007.06.05

The station is back online as of about day 156-16:20 after the station operator cycled power to the DAS.

2007.05.31

The station is down as of about day 151-19:20 due to communication problems between the DAS and the ISP.

2004.08.09

Station operating normally.

2003.09.13

Station operator replaced DAT drive, recording resumed 256-07:13 (9/13).

2003.09.08

Station operator reports DAT drive failure and no spare (9/4). Replacement shipped 9/4.

2003.01.17

New EPROMs installed to correct DSP problem.

2002.04.22

Station operating normally.